# Channel Dependence, Limited Lookback Windows, and the Simplicity of Datasets: How Biased is Time Series Forecasting?

### v2

@@ Appendix A Description of All Datasets
Second type of datasets used in our experiments is based on the chaotic ODE Benchmark. Here, we elaborate more on the general common features of the used ODE datasets and we give further description of the underlying processes for generating them. For the generation process, we use a common protocol for the chosen 6 attractors which is generating 20 points per time unit, splitting the data into 7:1:2 train:validation:test split which is the common practice for the standard datasets as well. Finally, we generate a very long time series 60000 timesteps which is divided into time series forecasting samples though a rolling window (same approach that is adopted on TSF papers for the standard datasets). For reproducibility, we provide the code stating the random seed that should be used to produce the same datasets and subsequently the same results we got. For understanding more the chosen processes as well as the dimensionality of each dataset, please refer to the chaotic ODE Benchmark paper (Gilpin, 2021).
Appendix B Hyperparameter Tuning
@@
We evaluate various Transformer-based and MLP-based time series forecasting models (e.g., DLinear, TSMixer, PatchTST, Crossformer, and our proposed FaCT) under a unified set of hyperparameter search ranges as shown in Table Table˜6.
@@
Hidden Dimension (d_modeld\_model) $\{128,256,512,1024\}$
Feedforward Dimension (d_ffd\_ff) $\{128,256,512,1024\}$
# Encoder/Mixer Layers [1, 10][1,\,10]
Dropout [0, 0.9][0,\,0.9]
@@
Table 6. Common Hyperparameter Search Ranges with all parameters being integer datatypes except for learning rate and dropout which span the whole range of floats on the respective range of values
For specific architectures, additional parameters that are model specific were tuned on the same range of parameters specified on the respective papers. For all the range $[]$ or set of values $\{\}$, refer to Table˜7.
@@
Crossformer/FaCT Segement Length $[3,12]$
@@
Table 7. Model Specific Hyperparameter Search Ranges/set of possible values with all of them having a datatype of integer in this case.
To ensure easy and full reproducibility of our experiments, we include all the best performing hyperparameter sets corresponding to all models and datasets. Please see tables from Table˜16 to Table˜28 for the full list. We hope inclusion of these lists, make reproducibility of this paper’s results straight forward.
Appendix C Implementation Details
C.1. Libraries and Hardware
All models were implemented using the PyTorch library (version 2.4) in Python (version 3.12.1). For hyperparameter optimization, we employed Optuna (version 3.6.1). Experiments were conducted on machines equipped with NVIDIA RTX 4090 and RTX 3090 GPUs.
C.2. Training Procedure
All models are trained using the Adam optimizer. We use mean squared error (MSE) as the default loss function, and each epoch iterates over mini-batches of size 32 by default. Training proceeds for a maximum of 100 epochs, or until early stopping is triggered if the validation loss fails to improve for 10 consecutive epochs. We use Optuna to conduct 20 trials per model, selecting configurations from the defined set of hyperparameters B that yield the best validation performance.
C.3. Granger Causality
Here, we list in more details the procedure to do the granger causality test which, alongside our code, should guide an easy usage of the framework for other datasets as well as easy reproduction of current results. We pick directly the first 1000 timesteps from the loaded dataset for this statistical test purposes, which we think is reasonable sample size with regard to the lags considered here. First, we check the pearson correlation between each pair of channels, and if it is too high (greater than 0,95 in our case), then one of the two channels is dropped. This is done, because channels which are very similar can cause results that are misleading when testing pair of channels that are identical (that can easily yield a very high score that will push the average scores way higher than expected). After, the initial filtering, we get dataset with diverse channels that can still have strong correlation, but one more important aspect we address next is stationarity as grangar causality works best on stationary time series data (i.e. mean and variance should not be shifting over time). Here, we apply the augmented Dickey-Fuller test on each channel of the time series to check if it is stationary. If the channel is not stationary, we apply differencing which should help bring stationarity into the time series channel. Now we have our preprocessed data which is ready for the grangar causality to be applied on. We use the packages from statsmodels package (Seabold and Perktold, 2010).
Appendix D Derivation of CI/CD version of Models
In our paper, 6 strong recent baselines are used throughout evaluation, they vary in complexity and structure. The aim of this appendix section is to ease the understanding of making CI and CD versions of the different baselines.
Simpler Linear/MLP-based Models
We start with the simpler models used in our experiments which are DLinear and TSMixer. DLinear as mentioned earlier is based on being the simplest model that applies a trend seasonality deceomposition followed by simpler linear layers applied on both the trend and the seasonal components (Zeng et al., 2023). For TSMixer, the authors had implemented a CD version which is the original version of the model and removed the channel mixing component which is called in their paper TMix-Only. We reimplemnted the model using pytorch and made sure of reproducing the paper results before using both versions in our experiments. Throughout our paper, we call both versions as TSMixer (CD) and TSMixer (CI) for the CD and CI versions respectively.
@@ Appendix D Derivation of CI/CD version of Models
Now we move into transformer-based models, starting with PatchTST as one of the first transformer models to re-establish a boost of performance over the DLinear model. In their paper, the original version of the model is CI where the transformer encoder backbone is applied independently on different channels before the application of a linear projection layer. For the CD version, they briefly discuss it in the appendix of applying the same backbone but on a flattened dimension of both number of patches and channels which means the attention is now applied jointly over patches and channels. We use the original implementation of the paper but we had to add the part related to CD implementation ourselves as it was not readily available on the paper at the time of writing our paper.
iTransformer/Crossformer
Next, transformer model that proved to be a strong baseline is Crossformer which applies a Depth-Segment-Wise embedding before application of a Two Stage Attention (TSA). Finally a decoder layer is applied through cross attention between the positional embedding of output time info and encoder output. The TSA is applied on the segments dimension (temporal) as well as followed by the channel dimension which constitutes the original CD version of the model (Zhang and Yan, 2023). For a CI version, we just removed the second stage of the TSA to make the attention only applied on the segments dimension. We used authors implementation of the CD version and made the mentioned edits on the implementation to produce the CI version used in our experiments. For iTransformer (Liu et al., 2024), we don’t introduce additional CI version to the existing CD version as otherwise the model will be simplified to a linear model which is already covered in this paper.
@@
One more prominent model proposed recently which proves be competitive on the standard datasets is TimeMixer (Wang et al., 2024a). In this model, the authors provide an MLP-based approach which is applied over multiple resolutions of the time series. This model original setting is CD where the MLP-mixing is done over both temporal and channel dimensions. For a CI version, the channel mixing component can be removed yielding a multiresolution temporal mixing based model. We use the authors implementation for both the CI and CD versions of the model.
Appendix E Detailed Results
In this appendix, we add the full results over all forecasting horizons for both types of introduced datasets. In Table˜9, we can see in detail the results for the chaotic ODE datasets, confirming the results from the main paper that Crossformer is clearly the SoTA on this benchmark with it being the best model in 15 out of possible 24 settings(for the CD version). For the standard datasets, the results are detailed in Table˜8 which also confirms the state of the current datasets where no model is clearly the best even emphasized with DLinear being the best model on 5 occurrences. This also fits the picture that was formulated on the table from the main paper text that when proper tuning is carried out for lookback window as well as all related hyperparameters, PatchTST remains competitive being the best overall 6 times. To conclude how comprehensive and reproducible our experiments are, these results represents the mean of 5 random seeds {3001, 3002, 3003, 3004, 3005} MSE error on the test split of each dataset.
Dataset / Model Horizon PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
@@ Appendix E Detailed Results
720 0.43* OOM 0.443 0.457 OOM 0.573* 0.466 OOM OOM 0.43*
Table 8. In this table, the full results over different horizons are shown over all the 7 standard datasets. These results were average of 5 different random seeds {3001,3002,3003,3004,3005}. The best result for each variant per model is in bold, while the best results overall for each horizon per dataset is highlighted in blue. we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
Dataset / Model Horizon PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
@@
Table 9. In this table, the full results over different horizons are shown over all the 6 chaotic ODE datasets. These results were average of 5 different random seeds {3001,3002,3003,3004,3005}. The best result for each variant per model is in bold, while the best results overall for each horizon per dataset is highlighted in blue. we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
Dataset / Model PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
@@
Traffic OOM OOM 0.0007 0.0025 0.0000 OOM OOM OOM OOM OOM
Table 10. In this table, all standard deviation values are averaged over the 4 used horizons for all datasets. For each horizon, the standard deviation results were acquired by running the model for 5 different seeds with the following values {3001,3002,3003,3004,3005}
. Note for some heavier models and datasets, we got Out of Memory error, specifically speaking some of the runs for electricity and traffic datasets.
Appendix F Detailed Granger Causality Results
We report results across three dataset groups (i) standard benchmark datasets, (ii) chaotic ODE datasets, and (iii) the large-scale Monash forecasting collection. For datasets with a large number of columns, we restrict the Granger causality analysis to pairwise comparisons among 20 randomly selected channels.
@@ Appendix F Detailed Granger Causality Results
For the standard benchmark datasets such as electricity and traffic, Granger causality tests reveal that these datasets exhibit moderate inter-channel correlation, with higher causality (F-score) predominantly at shorter lookback windows (lag=30) (Table˜11Table˜12, and Table˜14). This suggests that, for short lags, information from other channels can slightly aid forecasting, but as the lag increases, most predictive power lies within each channel’s own history. This outcome aligns with the observation that univariate models (CI) can often perform well for these datasets at longer lookback settings, as there is diminishing marginal benefit from including other channels when window size grows.
@@
In contrast, chaotic ODE datasets (Table˜11Table˜12, and Table˜14) consistently display higher Granger causality scores across all lag settings (with the exception of the electricity dataset), indicating robust and persistent dependencies between channels. This is particularly evident even at longer lookback windows. Such behavior is characteristic of deterministic physical systems or simulated multivariate dynamics where the evolution of one variable intrinsically influences the others. The implication is that multivariate models leveraging channel dependencies are especially suitable for these ODE datasets.
We also report results for the Monash repository’s (Godahewa et al., 2021) large-scale real-world datasets (over 40 time series from diverse domains). We select 40 datasets from a total of 58 datasets after removing datasets with missing values. Granger causality analysis Table˜13 reveals that only a minority—six of forty—show an inter-channel F-score above 10, and that F-scores tend to decline with increasing lag (except isolated cases like the Kaggle dataset). This pattern suggests that most Monash datasets lack strong lateral dependencies and instead behave like collections of independent univariate series. For instance, many resemble traffic datasets from the M1 or M3 collections: aggregations rather than integrated multivariate systems. As lags increase (e.g., at lag=196), additional channels yield minimal information gain, hence the F-score drops.
@@
Overall, these findings highlight that Granger causality excels at disentangling the nature of multivariate time series dependencies. Standard benchmarks typically permit effective univariate treatment at larger lags, the ODE datasets consistently demand modeling of channel interactions, and most Monash datasets demonstrate little true multivariate dependence. These results underscore the value of focusing research on genuinely multivariate datasets, such as chaotic ODEs, to advance understanding and modeling of channel dependence in time series forecasting.
Dataset / Metric Average F-score (30) $H_{0}$ rejected (%, p=0.05)
Double Pendulum 39.60 100%
Lorenz 1378.54 83%
Lorenz Coupled 579.23 97%
Cell Cycle 6.78 50%
Blinking Rotlet 3.98 33%
HopField 2.86 80%
ETTm1 1.48 50%
ETTm2 1.31 30%
ETTh1 1.87 60%
ETTh2 1.94 70%
Weather 1.55 33%
Electricity* (Ch=20) 4.34 93%
Traffic* (Ch=20) 4.71 94%
Table 11. Average F-scores and percentage of $H_{0}$ rejected (p=0.05) across datasets (lag=30).
Dataset / Metric Average F-score (96) $H_{0}$ rejected (%, p=0.05)
Double Pendulum 10.21 100%
Lorenz 353.25 83%
Lorenz Coupled 160.95 66%
Cell Cycle 2.61 53%
Blinking Rotlet 2.59 33%
HopField 1.92 70%
ETTm1 1.13 33%
ETTm2 1.11 15%
ETTh1 1.12 20%
ETTh2 1.19 36%
Weather 1.12 30%
Electricity* (Ch=20) 1.92 84%
Traffic* (Ch=20) 2.22 94%
Table 12. Average F-scores and percentage of $H_{0}$ rejected (p=0.05) across datasets (lag=96).
@@
australian electricity demand dataset 29.38 18.06 2.71 1.62
bitcoin without missing values 1.88 1.44 1.66 2.47
car parts without missing values 1.25
@@
kaggle web traffic without missing values 3.02 3.25 12.96 45.63
@@
kdd cup 2018 without missing values 2.48 1.53 1.18 1.08
london smart meters without missing values 1.69 1.26 1.07 1.03
m1 monthly 2.02
m1 quarterly
m1 yearly
m3 monthly 1.59
m3 other 1.46
m3 quarterly
m3 yearly
@@
m4 monthly 1.76
m4 quarterly 1.06
m4 weekly 1.14
m4 yearly
nn5 daily without missing values 9.45 2.10 1.17 1.02
@@
rideshare without missing values 3.73 2.15 1.31
@@
temperature rain without missing values 2.85 1.54 1.20 1.06
tourism monthly 12.12 1.58
tourism quarterly 1.98
tourism yearly
@@
vehicle trips without missing values 1.38
@@
Table 13. Average F-scores across 40 Monash datasets for lags 7, 30, 90, 192 .
Dataset / Metric Average F-score (192) $H_{0}$ rejected (%, p=0.05)
Double Pendulum 6.35 100%
Lorenz 126.38 83%
Lorenz Coupled 72.09 57%
Cell Cycle 1.58 53%
Blinking Rotlet 3.76 33%
HopField 1.55 50%
ETTm1 1.04 23%
ETTm2 0.99 0%
ETTh1 1.07 7%
ETTh2 1.06 17%
Weather 1.82 32%
Electricity* (Ch=20) 1.21 43%
Traffic* (Ch=20) 1.39 69%
Table 14. Average F-scores and percentage of rejected $H_{0}$ (p=0.05) across datasets (lag=192). Missing values indicate that the Granger Causality failed due to statistically limited number of available time points for a proper analysis.
Appendix G Full Lookback Window Analysis
We show here the detailed results of best performing lookback windows on seperate figures for each of the forecasting horizons. We can see from Figure˜4, Figure˜5, Figure˜6, and Figure˜7 the analysis of the results based on forecasting horizons 96, 192, 336, and 720 respectively. We can see that on average the same insights noticed from the discussed forecasting horizon 96, still applies to the rest of experiments. The patterns discussed with respect to ETTh1, and ETTh2 tending to use smaller lookback windows still applies clearly on horizons 192 and 336. This changes when 720 horizon is deployed which makes sense as the difficulty of forecasting such a large amount of timesteps promotes the usage of a larger lookback windows for all datasets.
Regarding the tendency of CI models to benefit more from a larger lookback window than CD models, that still holds for the different horizons as seen in the different figures. With also more models on the larger forecasting horizon, such as 720, tend to use larger lookback window more often. This pattern follows the same pattern we discussed for the performance on the ETTh datasets. The bottom line from this analysis is to show as well the complexity of forecasting different multivariate datasets across different horizons which makes the tuning of the lookback window clearly the standard approach to go for.

### v3

@@ Appendix 0.A Description of All Datasets
The second category of datasets is based on the Chaotic ODE Benchmark. We follow a common protocol for the six chosen attractors, generating 20 points per time unit and utilizing a 7:1:2 train:validation:test split. We generate a long time series of 60,000 timesteps, divided into forecasting samples via a rolling window. For reproducibility, we provide the code and random seeds required to generate these datasets. For further details on the dimensionality and underlying processes, we refer to the Chaotic ODE Benchmark paper [5].
Appendix 0.B Hyperparameter Tuning
@@
We evaluate various Transformer-based and MLP-based models under a unified hyperparameter search space (Table 6).
Table 6: Common Hyperparameter Search Ranges. Values are integers except for learning rate and dropout (floats).
@@
Hidden Dimension ($d_{model}$) $\{128,256,512,1024\}$
Feedforward Dimension ($d_{ff}$) $\{128,256,512,1024\}$
# Encoder/Mixer Layers $[1,10]$
Dropout $[0,0.9]$
@@
Model-specific parameters were tuned within the ranges suggested by their respective original papers (Table 7). Best-performing hyperparameter sets for each model-dataset pair are included in the supplementary materials to ensure full reproducibility.
Table 7: Model Specific Hyperparameter Search Ranges (all values are integers).
@@
Crossformer/FaCT Segment Length $[3,12]$
@@
Appendix 0.C Implementation Details
0.C.1 Libraries and Hardware
All models were implemented using PyTorch (v2.4) in Python (v3.12.1). Hyperparameter optimization was conducted via Optuna (v3.6.1). Experiments were performed on machines equipped with NVIDIA RTX 4090 and RTX 3090 GPUs.
0.C.2 Training Procedure
Models were trained using the Adam optimizer with Mean Squared Error (MSE) loss. The default mini-batch size was 32. Training lasted for a maximum of 100 epochs, with early stopping triggered if validation loss did not improve for 10 consecutive epochs.
0.C.3 Granger Causality
We utilize a Granger causality framework to analyze inter-channel dependencies. We extract the first 1,000 timesteps for this statistical test. Initially, we filter out highly redundant channels (Pearson correlation >0.95>0.95). We then apply the Augmented Dickey-Fuller (ADF) test to ensure stationarity, applying differencing where necessary. The causality tests are performed using the statsmodels package [20].
Appendix 0.D Derivation of CI/CD Model Variants
We evaluated six recent baselines, creating both Channel-Independent (CI) and Channel-Dependent (CD) versions where applicable.
Simpler Linear/MLP Models
DLinear remains a CI baseline due to its decomposition approach. For TSMixer, we implemented a CI version by removing the channel-mixing component (referred to as TMix-Only in the original work). Both versions are denoted as TSMixer (CI) and TSMixer (CD).
@@ Appendix 0.D Derivation of CI/CD Model Variants
The original PatchTST is CI. For our CD variant, we apply the transformer backbone to a flattened dimension of both patches and channels, enabling joint attention across both.
Crossformer
The standard CD version uses Two-Stage Attention (TSA) over temporal segments and channels. Our CI version removes the second stage of TSA to restrict attention to the temporal dimension only. For iTransformer, we utilize only the CD version, as a CI version would simplify the model into a basic linear projection already covered by other baselines.
@@
Originally a CD model utilizing multi-resolution MLP-mixing. We derived the CI version by removing the channel-mixing components while maintaining the multi-resolution temporal mixing structure.
Appendix 0.E Detailed Results
Full forecasting results across all horizons are presented in Table 9 (ODE datasets) and Table 8 (Standard datasets). Results represent the mean MSE over five random seeds: $\{3001,3002,3003,3004,3005\}$.
Table 8: Full forecasting results (MSE) over standard datasets. Best result per model variant in bold; best overall result highlighted in blue.
PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
Dataset H CI CD CI CD CI CD CI CD CI CD
@@ Appendix 0.E Detailed Results
720 0.430* OOM 0.443 0.457 OOM 0.573* 0.466 OOM OOM 0.430*
Table 9: Full forecasting results (MSE) for Chaotic ODE datasets.
PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
Dataset H CI CD CI CD CI CD CI CD CI CD
@@
Table 10: Standard deviation values averaged over the 4 used horizons. "—" denotes Out of Memory (OOM) errors.
PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
Dataset CI CD CI CD CI CD CI CD CI CD
@@
Traffic 0.0007 0.0025 0.0000
Appendix 0.F Detailed Granger Causality Analysis
We evaluate inter-channel dependencies across three distinct dataset groups:
  1. (i)
    Standard benchmark datasets (e.g., ETT, Weather);
  2. (ii)
    Chaotic ODE datasets (e.g., Lorenz, Double Pendulum);
  3. (iii)
    The large-scale Monash forecasting collection [6].
For datasets with high dimensionality, we restrict the Granger causality analysis to pairwise comparisons among 20 randomly selected channels to maintain computational efficiency.
Standard Benchmarks
@@ Appendix 0.F Detailed Granger Causality Analysis
As shown in Table 11, datasets such as Electricity and Traffic exhibit moderate inter-channel correlation. However, higher causality ($F$-scores) is predominantly concentrated at shorter lookback windows (lag=30). As the lag increases to 192, the predictive power of neighboring channels diminishes significantly. This quantitative decline supports the observation that Channel-Independent (CI) univariate models often perform optimally at longer lookback settings, as the marginal information gain from other channels becomes negligible over time.
Chaotic ODE Datasets
@@
In contrast, chaotic ODE datasets consistently display high $F$-scores across all lag settings. Even at a lag of 192, systems like the Lorenz and Double Pendulum maintain robust and persistent dependencies between channels. This is characteristic of deterministic physical systems where the state of one variable is intrinsically and permanently linked to the evolution of others. These results underscore why multivariate models (CD) are particularly well-suited for physical or simulated dynamics.
Monash Forecasting Collection
We analyzed 40 datasets from the Monash repository after filtering for missing values. As detailed in Table 12, the vast majority of these real-world datasets lack strong lateral dependencies. Only six of the forty datasets show an inter-channel $F$-score above 10, with most scores declining sharply as lag increases. Many of these datasets (e.g., Hospital or Australian Electricity) behave more like collections of independent univariate series rather than integrated multivariate systems.
Summary
@@
Overall, Granger causality serves as a powerful diagnostic tool for disentangling the nature of time series dependencies. While standard benchmarks permit effective univariate treatment at larger lags, ODE datasets demand explicit modeling of channel interactions. These findings suggest that research into multivariate dependencies should prioritize genuinely integrated datasets, such as chaotic ODEs, rather than simple aggregations.
Table 11: Average $F$-scores and percentage of $F$-tests passed ($p=0.05$) across different lag settings. Asterisks (*) indicate analysis performed on 20 randomly selected channels.
Lag = 30 Lag = 96 Lag = 192
Dataset F-score Passed % F-score Passed % F-score Passed %
Double Pendulum 39.60 100% 10.21 100% 6.35 100%
Lorenz 1378.54 83% 353.25 83% 126.38 83%
Lorenz Coupled 579.23 97% 160.95 66% 72.09 57%
Cell Cycle 6.78 50% 2.61 53% 1.58 53%
Blinking Rotlet 3.98 33% 2.59 33% 3.76 33%
HopField 2.86 80% 1.92 70% 1.55 50%
ETTm1 1.48 50% 1.13 33% 1.04 23%
ETTm2 1.31 30% 1.11 15% 0.99 0%
ETTh1 1.87 60% 1.12 20% 1.07 7%
ETTh2 1.94 70% 1.19 36% 1.06 17%
Weather 1.55 33% 1.12 30% 1.82 32%
Electricity* 4.34 93% 1.92 84% 1.21 43%
Traffic* 4.71 94% 2.22 94% 1.39 69%
Table 12: Average $F$-scores across 40 Monash datasets for lags 7, 30, 96, and 192. Empty cells or "—" indicate insufficient time points for statistical validity at that lag.
@@
Australian electricity demand 29.38 18.06 2.71 1.62
Bitcoin (no missing) 1.88 1.44 1.66 2.47
Car parts (no missing) 1.25
@@
Kaggle web traffic (no missing) 3.02 3.25 12.96 45.63
@@
KDD Cup 2018 (no missing) 2.48 1.53 1.18 1.08
London smart meters (no missing) 1.69 1.26 1.07 1.03
M1 Monthly / Quarterly / Yearly 2.02
M3 Monthly / Other / Quarterly / Yearly 1.59
@@
M4 Monthly / Quarterly / Weekly 1.76
NN5 Daily (no missing) 9.45 2.10 1.17 1.02
@@
Rideshare (no missing) 3.73 2.15 1.31
@@
Temperature rain (no missing) 2.85 1.54 1.20 1.06
Tourism Monthly / Quarterly / Yearly 12.12 1.58
@@
Vehicle trips (no missing) 1.38
@@
Appendix 0.G Full Lookback Window Analysis
This section provides a granular analysis of optimal lookback window frequencies across all evaluated forecasting horizons ($H\in\{96,192,336,720\}$). Figures 4 through 7 detail these results, categorized by dataset, Channel-Independent (CI) models, and Channel-Dependent (CD) models.
On average, the insights derived from the $H=96$ horizon remain consistent across the extended experiments. For instance, the tendency of ETTh1 and ETTh2 to utilize smaller lookback windows persists for horizons 192 and 336. However, at $H=720$, we observe a systemic shift toward larger lookback windows. This transition is intuitive; the increased difficulty of forecasting across an extended horizon promotes the use of broader historical context for all datasets.
Furthermore, the observation that CI models typically benefit more from larger lookback windows than CD models holds true across horizons. As the forecasting horizon expands to 720, a greater number of models across all categories favor larger windows. These patterns underscore the complexity of multivariate time series forecasting and confirm that the systematic tuning of lookback windows is essential for optimal performance across diverse horizons.
(a) By Dataset
(b) By CI Models
(c) By CD Models